OpenStudy (avai70178):

Someone please help. I understand the chain rule for derivatives but i'm not sure how to use it in this problem. y=(x^2+1)(x+1)^2

OpenStudy (loser66):

It is not a chain rule, it is product rule. Where are you stuck?

OpenStudy (avai70178):

I'm not sure if I should use the product rule on both pieces of the equation with the same exponent.

OpenStudy (loser66):

Let a = x^2 +1 then a' = ? Let b = (x+1)^2 , then b' =?

OpenStudy (avai70178):

a= 2 b=2x+2

OpenStudy (loser66):

no, not that a' = 2x, b' is ok.

OpenStudy (loser66):

ok, now. y = a*b , then y' = a'b + b'a ok?

OpenStudy (avai70178):

oops sorry i didnt mean to post that

OpenStudy (avai70178):

wait well what is a then?

OpenStudy (avai70178):

if we are using the product rule than a should equal just 2

OpenStudy (loser66):

is it not that we defined a = x^2 +1? and b = (x+1)^2??

OpenStudy (avai70178):

yes and if we take the derivatives of those then we have a'=2 and b'=?

OpenStudy (loser66):

I think you got confuse. ok, let me walk you through the concept. Answer me one by one ok? (x) ' =?

OpenStudy (avai70178):

do you mean derivative of x? cause idk

OpenStudy (loser66):

yes, derivative of x =?

OpenStudy (avai70178):

idk

OpenStudy (avai70178):

im sorry i feel dumb

OpenStudy (loser66):

ok, I tell you x' =1 \((x^\color{red}{2})' =\color{red}{2}x^1= \color{red}{2}x\) \((x^\color{red}{3})' =\color{red}{3}x^2\) \((x^\color{red}{5})' =\color{red}{5}x^4\) Now, tell me \((x^{10})' =?\)

OpenStudy (avai70178):

10x^9

OpenStudy (loser66):

Good, that is called "n-rule" , ok?

OpenStudy (avai70178):

ok

OpenStudy (loser66):

One more thing you should know, derivative of a constant =0, that is (3)' =0 5' =0 10'=0 tell me (234)' =?

OpenStudy (avai70178):

0

OpenStudy (loser66):

good!! Now, combine them \((x +y) ' = x' +y' \) \((x^2 +5)' = (x^2)' +5' =???\)

OpenStudy (avai70178):

2x?

OpenStudy (loser66):

yup

OpenStudy (loser66):

now, product rule: (x*y)' = x'*y + y' *x ok?

OpenStudy (avai70178):

yes

OpenStudy (loser66):

Now, Your problem : \(((x^2+1) *(x+1)^2)' \)

OpenStudy (loser66):

apply product rule, right? what do you get?

OpenStudy (avai70178):

8x+4

OpenStudy (avai70178):

i think